# Towards Generalisable Time Series Understanding Across Domains

### v1

@@ 4.4 Domain Signature Analysis
A key component of OTiS is its use of domain-specific variate embeddings. While these embeddings are randomly initialised, we expect them to capture unique domain characteristics during training, eventually serving as the signature of their respective domain. To validate this hypothesis, we analyse the domain-specific variate embeddings after pre-training using principal component analysis (PCA).
First, we find that OTiS unifies time series from diverse domains into a meaningful latent space, where embeddings of domains with shared high-level semantics cluster together, as depicted in Appendix E.1. For example, embeddings of mono and stereo audio group closely, as do those of banking and economics. Moreover, EEG-specific embeddings are clearly separated and ECG-specific embeddings form a tight cluster.
Second, OTiS preserves low-level semantics specific to each domain, such as the relationships between variates. To explore this, we focus on the learned variate embeddings of EEG, the most complex domain in our corpus. EEG variates correspond to actual electrodes, each associated with a $3$D position in space or a $2$D position on the scalp, which is ideal for studying inter-variate relationships. Our analysis covers both (i) variate embeddings for $10$-$20$ system EEG recordings with $19$ electrodes learned during multi-domain pre-training, and (ii) variate embeddings for previously unseen EEG recordings with $32$ electrodes learned during fine-tuning of the pre-trained OTiS. We find that the first three principal components explain (i) $74.7\,\%$ and (ii) $87.9\,\%$ of the variance, indicating that the $3$D PCA projections of the variate embeddings correspond closely to the actual electrode positions. A visualisation of the learned $3$D electrode layout is provided in Appendix E. Additionally, we linearly align these positions with the true EEG electrode layout on the $2$D scalp manifold, as depicted in Figure 3. The alignment yields $R^{2}$ values of (i) $0.81$ and (ii) $0.95$, confirming a strong correspondence between the learned variate embeddings and the actual electrode layout.
Figure 3: $2$D PCA projections of the learned EEG-specific variate embeddings, overlaid on the actual EEG electrode layout. (Left) Embeddings for $10$-$20$ system EEG recordings with $19$ electrodes learned during pre-training. (Right) Embeddings for previously unseen EEG recordings with $32$ electrodes learned during fine-tuning. The embeddings accurately reflect the spatial electrode layout, as confirmed by high correlations (R2) between the PCA projections $\bullet$ and the true layout $\circ$.
@@ 4.5 Scaling Study
Figure 4: Performance of OTiS with different numbers of pre-training samples. Shaded regions indicate the standard deviation across $5$ seeds. Increasing dataset size generally improves downstream performance. Scaling model size requires even larger pre-training corpora to be effective.
We analyse the scaling behaviour of OTiS with respect to model and dataset size. To this end, we subsample the pre-training data to $10\,\%$ and $1\,\%$ of its original size, ensuring that each subset is fully contained within the corresponding superset. We evaluate the downstream performance of all OTiS variants across classification, regression, and forecasting tasks, as depicted in Figure 4.
The experiments demonstrate that downstream performance generally scales with dataset size, achieving the best results with the full pre-training dataset. This trend, however, does not directly apply to model size, which is in line with the scaling behaviour observed in current time series foundational models (Woo et al., 2024; Goswami et al., 2024). Given that performance generally improves across all models with increasing data size, we hypothesise that scaling the model size could prove beneficial with even larger pre-training corpora.
@@ 4.6 Ablation Study
Figure 5: Ablation study on key components of OTiS. Downstream performance is analysed across $5$ seeds. A leave-one-out approach is used to evaluate the influence of each component. The default setting, which includes all components, demonstrates superior model capabilities across tasks.
We perform an ablation study to analyse the impact of OTiS’ key components: the domain-specific tokeniser, dual masking strategy, and normalised cross-correlation (NCC) loss. As shown in Figure 5, the best and most robust performance is achieved when all components are used during pre-training.
Replacing the domain-specific variate embeddings with domain-agnostic embeddings (i.e. learnable embeddings shared across all domains) consistently led to inferior performance across all tasks, demonstrating the importance of capturing domain-specific data characteristics during tokenisation. Switching from dual masking to random masking resulted in performance degradation, although the impact was less notable for generative tasks than for discriminative tasks. We hypothesise that the NCC loss already captures temporal causality, which is particularly crucial for generative tasks like forecasting. Overall, removing the NCC loss caused performance declines across all downstream tasks, emphasising the role of long-range dependencies for general time series understanding.
@@ 5 Discussion & Conclusion
In this study, we explore the problem of effective pre-training on heterogeneous time series corpora. Time series vary substantially across domains, e.g. with respect to inter-variate relationships and temporal dynamics, rendering generalisable feature extraction from multi-domain time series difficult. To address this issue, we present OTiS, an open model for general time series analysis, specifically designed to handle multi-domain heterogeneity. Our novel multi-domain pre-training paradigm, including a domain-specific tokeniser with learnable signatures, a dual masking strategy, and a normalised cross-correlation (NCC) loss, enables OTiS to extract generalisable time series features.
In extensive experiments, we demonstrate that OTiS generalises well across $15$ diverse downstream applications spanning $8$ distinct domains, achieving competitive performance with both specialised and general state-of-the-art (SOTA) models. In a qualitative analysis, we further show that OTiS unifies time series from diverse domains in a meaningful latent space, while preserving low-level semantics of a domain including the inter-variate relationships. Thereby, our work establishes a strong foundation for future advancements in interpretable and general time series analysis.
@@
While OTiS outperforms SOTA models across $10$ tasks, our experiments in low-data regimes suggest that larger pre-training corpora could further enhance its performance. Unlike in NLP and CV, where large datasets are curated from web-crawled data, foundational models in time series, including OTiS, still rely on manually curated datasets. Future work could explore fully automatic pipelines, e.g. using embedding similarity, to filter and rebalance multi-domain time series from the web. OTiS could further benefit from processing domain signatures during inference, potentially unlocking zero-shot capabilities, similarly to those seen in foundational models in NLP and CV.
@@ Appendix A Large Multi-Domain Pre-Training Corpus
The AudioSet dataset (Gemmeke et al., 2017) contains $10$-second YouTube clips for audio classification, featuring $527$ types of audio events that are weakly annotated for each clip. The full training set includes a class-wise balanced subset (AudioSet-$20$K, $22,176$ clips) and an unbalanced (AudioSet-$2$M $2,042,985$ clips) set. For our pre-training corpus, we use the balanced AudioSet-$20$K, which contains $3,491$ mono and $16,123$ stereo recordings, all sampled at $44,100\,$Hz.
@@
Appendix B Benchmark Dataset Details
We provide an overview of the datasets used to benchmark our model on $15$ distinct downstream applications in Table 4. Our benchmarking experiments include classifcation, regression, and forecasting tasks.
Table 4: Experimental setup with all conducted downstream applications, evaluation metrics, and datasets used for benchmarking.
@@ Appendix B Benchmark Dataset Details
Forecasting MSE Energy ETTh1 2021 $1$ $7$ $17,420$ (hourly) 278278278\,278mHz
ETTh2 2021 $1$ $7$ $17,420$ (hourly) 278278278\,278mHz
@@
Electricity Electricity 2024 $321$ $1$ $26,304$ (hourly) 278278278\,278mHz
@@ C.2 Pre-Training & Fine-Tuning Parameters
We provide the hyperparameters used to pre-train all variants of OTiS in Table 6. The hyperparameters used to fine-tune our models for the classification, regression, and forecasting tasks are provided in Table 7, 8, and 9, respectively.
@@
Gesture OTiS-Base $75$ $32$ $1$e-$2$ $0.1$ $0.25$ $0.2$ $0.1$
@@
ETTh1 OTiS-Base $1000$ $1$ $1$e-00 $0.1$ $0.15$
OTiS-Large $1000$ $1$ $1$e-$1$ $0.2$ $0.15$
OTiS-Huge $1000$ $1$ $3$e-$1$ $0.1$ $0.15$
ETTh2 OTiS-Base $1000$ $1$ $1$e-00 $0.2$ $0.25$
OTiS-Large $1000$ $1$ $1$e-$1$ $0.1$ $0.25$
OTiS-Huge $1000$ $1$ $3$e-$1$ $0.0$ $0.25$
ETTm1 OTiS-Base $1000$ $1$ $3$e-$1$ $0.2$ $0.25$
OTiS-Large $1000$ $1$ $3$e-$1$ $0.2$ $0.25$
OTiS-Huge $1000$ $1$ $1$e-$1$ $0.1$ $0.15$
ETTm2 OTiS-Base $1000$ $1$ $3$e-$1$ $0.1$ $0.25$
OTiS-Large $1000$ $1$ $3$e-$1$ $0.2$ $0.25$
OTiS-Huge $1000$ $1$ $1$e-$1$ $0.2$ $0.25$
Weather OTiS-Base $1000$ $1$ $3$e-$1$ $0.2$ $0.15$
OTiS-Large $1000$ $1$ $3$e-$1$ $0.2$ $0.15$
OTiS-Huge $1000$ $1$ $1$e-$1$ $0.2$ $0.05$
Electricity OTiS-Base $250$ $32$ $3$e-$2$ $0.0$ $0.25$
OTiS-Large $250$ $32$ $3$e-$2$ $0.0$ $0.15$
OTiS-Huge $250$ $32$ $3$e-$2$ $0.2$ $0.15$
@@ Appendix E Domain Signature Analysis
To analyse the domain signatures, we reduce the dimensionality of the domain-specific variate embeddings by employing a principal component analysis (PCA). Our analysis shows that OTiS unifies time series from diverse domains into a meaningful latent space, while accurately capturing the inter-variate relationships within a domain.
@@ E.1 Inter-Domain Analysis
A visualisation of all domain-specific variate embeddings learned during pre-training is provided in Figure 6. We find that OTiS learns a meaningful latent space, where embeddings of domains with shared high-level semantics cluster closely together.
@@ E.2 Intra-Domain Analysis
A visualisation of the domain-specific variate embeddings for EEG and ECG are given in Figure 7 and Figure 8, respectively. We find that OTiS preserves low-level semantics specific to each domain, accurately capturing the relationships between variates.
(a) True electrode layout.
(b) Implicitly learned electrode layout.
@@
Figure 7: $3$D PCA projections of the variate embeddings for $10$-$20$ system EEG recordings with $19$ electrodes learned during pre-training. The labels of each node correspond to the electrode names.

### v2

@@ 4.5 Domain Signature Analysis
Inter-Variate Relationships.
A key design choice of OTiS is the use of domain-specific variate embeddings. While these embeddings are randomly initialised, we expect them to capture unique domain characteristics, e.g. inter-variate relationships, during training. To test this hypothesis, we analyse the EEG-specific and ECG-specific variate embeddings learned during pre-training, and the Weather-specific variate embeddings learned during fine-tuning.
The principal component analysis (PCA) presented in Figure 7 (see Appendix E.2.1) demonstrates that EEG-specific variate embeddings accurately capture the $3$D spatial arrangement of EEG electrodes placed on the human scalp. Similarly, the PCA in Figure 9 (see Appendix E.2.1) indicates that ECG-specific variate embeddings correctly reflect the relationships among the $12$ leads: limb leads I, II, and III form a planar $2$D triangle, representing Einthoven’s triangle (Kligfield et al., 2007), while precordial leads V1111–V6666 capture a $3$D view of the heart. Moreover, the embedding similarity analysis in Figure 10 (see Appendix E.2.1) reveals that Weather-specific variate embeddings capture the true physical relationships among climatological indicators. This also showcases that OTiS successfully adapts to previously unseen domains with limited data seen during fine-tuning.
Temporal Patterns.
The central question is whether OTiS learns only the relationships across variates or also captures temporal patterns within a variate. To investigate this, we conduct forecasting experiments on uni-variate sine waves with distinct frequencies. Note that in a uni-variate setting ($V_{\text{sine}}=1$), we ensure that our model does not utilise information from other variates. For these experiments, we freeze both the encoder $f(\cdot)$ and decoder $g(\cdot)$ after pre-training and fine-tune only a randomly initialised variate embedding $\bm{E}^{\mathcal{V}}_{\text{sine}}\in\mathbb{R}^{1\times D}$ ($<0.2\,$k trainable parameters) on 505050\,50Hz sine waves. During inference, OTiS is fully frozen and forecasts are made on unseen sine waves with frequencies of 222\,2Hz, 282828\,28Hz, 606060\,60Hz, and 100100100\,100Hz. As shown in Figure 11 (see Appendix E.2.2), we find that the domain-specific variate embedding $\bm{E}^{\mathcal{V}}_{\text{sine}}$ captures temporal patterns that generalise to previously unseen frequencies. These observations demonstrate that OTiS learns not only the inter-variate relationships of a domain, but also its temporal patterns.
@@ 4.6 Scaling Study
Figure 4: Scaling study. Shaded regions indicate the standard deviation across $5$ seeds. Downstream performance across tasks generally scales with dataset size. Scaling the model size requires even larger pre-training corpora to be effective.
We analyse the scaling behaviour of OTiS with respect to dataset and model size. To this end, we subsample the pre-training data to $10\,\%$ and $1\,\%$ of its original size, ensuring that each subset is fully contained within the corresponding superset. We evaluate the downstream performance of all OTiS variants across classification, regression, and forecasting tasks, as depicted in Figure 4.
The experiments demonstrate that downstream performance generally scales with dataset size, achieving the best results with the full pre-training dataset. This trend, however, does not directly apply to model size, which is in line with the scaling behaviour observed in current time series foundation models (Woo et al., 2024; Goswami et al., 2024). Given that an increasing dataset size generally improves model performance, we hypothesise that scaling the model size could prove beneficial with even larger pre-training corpora.
@@ 4.7 Ablation Study
We conduct an ablation study to analyse the impact of OTiS’ key components: the domain-specific tokeniser, dual masking strategy, and normalised cross-correlation (NCC) loss. As shown in Figure 5, OTiS achieves superior performance when all components are integrated during pre-training.
Replacing domain-specific variate embeddings with domain-agnostic ones (i.e. variate embeddings shared across all domains) consistently led to inferior performance across tasks, indicating the importance of considering domain characteristics during tokenisation. Switching from dual to random masking resulted in performance degradation, although the impact was less notable for generative tasks than for discriminative tasks. We hypothesise that the NCC loss already captures temporal causality, which is particularly helpful for generative tasks like forecasting. Overall, removing the NCC loss caused performance declines across all downstream applications, emphasising the importance of long-range dependencies for general time series understanding.
Figure 5: Ablation study. Downstream performance is analysed across $5$ seeds. A leave-one-out approach is used to evaluate the influence of each component. The default setup, incorporating all components, achieves superior performance across tasks.
@@ 5 Discussion & Conclusion
In this study, we explore the problem of efficient pre-training on heterogeneous time series corpora. Prevalent in domains ranging from medicine to finance, time series vary notably e.g with respect to inter-variate relationships and temporal patterns, rendering efficient pre-training across domains difficult. To address this, we propose a novel pre-training paradigm including a tokeniser with learnable domain signatures, a dual masking strategy, and a normalised cross-correlation loss, enabling our open model for general time series analysis (OTiS) to learn general time series features.
In extensive benchmarking experiments, OTiS is competitive against specialised state-of-the-art (SoTA) models. Our qualitative analysis highlights how learnable domain signatures address time series heterogeneity, facilitating the disentanglement of domain-specific knowledge and enabling OTiS to acquire knowledge on fundamental time series properties shared across domains, such as frequency, amplitude, offset, or phase. This is particularly valuable in domains with limited data availability and high privacy standards, such as medicine, where our approach offers a promising solution by leveraging readily available data from any domain. Overall, our work establishes an essential foundation for future advancements in time series analysis.
@@
While OTiS outperforms SoTA models across diverse tasks, scaling studies indicate that larger pre-training corpora could further benefit performance. Unlike in NLP and CV, where large corpora are curated from the web, time series foundation models, including OTiS, still rely on manually curated data. Future work could explore fully automated pipelines to filter time series from the web and categorise them into domains, e.g. using embedding similarities. Finally, integrating other modalities, such as imaging or text, is a promising direction for future research.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Large Multi-Domain Pre-Training Corpus
The AudioSet dataset (Gemmeke et al., 2017) contains $10$-second YouTube clips for audio classification, featuring $527$ types of audio events that are weakly annotated for each clip. The full training set includes a class-wise balanced subset (AudioSet-$20$K, $22,176$ clips) and an unbalanced (AudioSet-$2$M $2,042,985$ clips) set. For our pre-training corpus, we use the balanced AudioSet-$20$K, which contains $3,491$ mono and $16,123$ stereo recordings, all sampled at 44.144.144.1\,44.1kHz.
@@
Appendix B Benchmark Details
To assess the utility of OTiS in real-world settings, we conduct experiments on three key use cases in time series analysis: classification, regression, and forecasting. For classification, we perform binary epilepsy detection using EEG (Epilepsy 2001), multi-class fault detection in rolling bearings from vibration signals (FD-B 2016), multi-class hand-gesture classification with accelerometer signals (Gesture 2009), and multi-class muscular disease classification using electromyographie (EMG 2000). For regression, we predict five imaging-derived cardiac phenotypes from $12$-lead ECG (LVEDV, LVESV, LVSV, LVEF, LVM 2020). For forecasting, we predict electricity transformer temperature (ETT 2021), weather (Weather 2024), and electricity consumption (Electricity 2024). We adhere to the established data splitting and evaluation procedures for the classification (Zhang et al., 2022), regression (Turgut et al., 2025), and forecasting (Zhou et al., 2021) tasks. We provide an overview of the datasets and baselines used to benchmark our model in Table 4 and Table 5, respectively.
Table 4: Summary of all datasets used for benchmarking, including evaluation metrics, domains, and dataset details.
@@ Appendix B Benchmark Details
Forecasting MSE Energy ETTh1 2021 $1$ $7$ $17,420$ (hourly) $278\,\mu$Hz
ETTh2 2021 $1$ $7$ $17,420$ (hourly) $278\,\mu$Hz
@@
Electricity Electricity 2024 $321$ $1$ $26,304$ (hourly) $278\,\mu$Hz
Table 5: Summary of all baseline models used for benchmarking, including pre-training details, domain adaptation methods, and architectural design. CL, MDM, and GPT denote contrastive learning, masked data modelling, and generative pre-training, respectively.
Task Model Pre-training Domain adaptation Architecture
Method Dataset
Classification SimCLR 2020 CL SleepEEG* 2000 Fine-tuning $1$D-CNN
TimesNet 2022a Fine-tuning $2$D-CNN
CoST 2022 CL SleepEEG* 2000 Fine-tuning $1$D-CNN
TS2Vec 2022 CL SleepEEG* 2000 Fine-tuning $1$D-CNN
TF-C 2022 CL SleepEEG* 2000 Fine-tuning Transformer
Ti-MAE 2023 MDM SleepEEG* 2000 Fine-tuning Transformer
SimMTM 2024 MDM SleepEEG* 2000 Fine-tuning Transformer
ST-Transformer 2021 Fine-tuning Transformer
CNN-Transformer 2022 Fine-tuning CNN w/ Transformer
FFCL 2022 Fine-tuning CNN w/ LSTM
SPaRCNet 2023 Fine-tuning $1$D-CNN
ContraWR 2023 CL Target Fine-tuning Transformer
BIOT 2024 MDM + Fine-tuning Transformer
LaBraM 2024 MDM Fine-tuning Transformer
Regression iTransformer 2023 Fine-tuning Transformer
ViT 2025 Fine-tuning Transformer
MAE 2025 MDM Target Fine-tuning Transformer
CM-AE 2023 MDM and CL Target Fine-tuning $1$D-CNN
MMCL 2025 MDM and CL Target Fine-tuning Transformer
Forecasting N-BEATS 2019 Fine-tuning Non-Linear Model
Autoformer 2021 Fine-tuning Transformer
TimesNet 2022a Fine-tuning $2$D-CNN
DLinear 2023 Fine-tuning Linear Model
PatchTST 2023 MDM Target Fine-tuning Transformer
Time-LLM 2023 GPT Fine-tuning Transformer
GPT4TS 2023 GPT Fine-tuning Transformer
MOMENT 2024 MDM TSP 2024 Fine-tuning Transformer
MOIRAI 2024 MDM LOTSA 2024 Zero-shot Transformer
* $371,055$ uni-variate, 2-seconds EEG recordings sampled at a frequency of $100$ Hz.
+ Pre-training corpus spans $6$ EEG datasets with more than $5$ million samples and $13,000$ recording hours
Pre-training corpus spans $16$ EEG datasets with a total of $2,500$ recording hours
Model uses Llama-$7$B 2023, pre-trained on $1.4$ trillion text tokens, as encoder.
Model uses GPT2222 2018, pre-trained on $10$ billion text tokens, as encoder.
Time Series Pile (TSP) contains $13$ million samples and $1.23$ billion time points from $13$ domains.
Large-Scale Open Time Series Archive (LOTSA) contains more than $4$ million samples and $27$ billion time points from $9$ domains.
@@ C.2 Pre-Training & Fine-Tuning Parameters
We provide the hyperparameters used to pre-train all variants of OTiS in Table 7. The hyperparameters used to fine-tune our models for the classification, regression, and forecasting tasks are provided in Table 8, 9, and 10, respectively.
@@
Gesture OTiS-Base $75$ $32$ $3$e-$3$ $0.2$ $0.50$ $0.1$ $0.1$
@@
ETTh1 OTiS-Base $1000$ $1$ $3$e-$2$ $0.1$ $0.15$
OTiS-Large $1000$ $1$ $3$e-$3$ $0.2$ $0.15$
OTiS-Huge $1000$ $1$ $1$e-$2$ $0.1$ $0.15$
ETTh2 OTiS-Base $1000$ $1$ $3$e-$2$ $0.2$ $0.25$
OTiS-Large $1000$ $1$ $3$e-$3$ $0.1$ $0.25$
OTiS-Huge $1000$ $1$ $1$e-$2$ $0.0$ $0.25$
ETTm1 OTiS-Base $1000$ $1$ $1$e-$2$ $0.2$ $0.25$
OTiS-Large $1000$ $1$ $1$e-$2$ $0.2$ $0.25$
OTiS-Huge $1000$ $1$ $3$e-$3$ $0.1$ $0.15$
ETTm2 OTiS-Base $1000$ $1$ $1$e-$2$ $0.1$ $0.25$
OTiS-Large $1000$ $1$ $3$e-$3$ $0.2$ $0.25$
OTiS-Huge $1000$ $1$ $1$e-$2$ $0.2$ $0.25$
Weather OTiS-Base $1000$ $1$ $1$e-$2$ $0.2$ $0.25$
OTiS-Large $1000$ $1$ $1$e-$2$ $0.2$ $0.15$
OTiS-Huge $1000$ $1$ $3$e-$3$ $0.2$ $0.05$
Electricity OTiS-Base $250$ $32$ $1$e-$3$ $0.0$ $0.25$
OTiS-Large $250$ $32$ $1$e-$3$ $0.0$ $0.15$
OTiS-Huge $250$ $32$ $1$e-$3$ $0.2$ $0.15$
@@ Appendix E Domain Signature Analysis
We analyse the domain signatures, i.e. the domain-specific variate embeddings, by employing a principal component analysis (PCA) and embedding similarity analysis. Our analyses show that OTiS unifies time series from diverse domains into a meaningful latent space (see Section E.1), while accurately capturing the inter-variate relationships (see Section E.2.1) and temporal patterns within a domain (see Section E.2.2).
@@ E.1 Inter-Domain Analysis
A visualisation of all domain-specific variate embeddings learned during pre-training is provided in Figure 6. OTiS learns a meaningful latent space, where embeddings of domains with shared high-level semantics cluster closely together.
@@ E.2 Intra-Domain Analysis
E.2.1 Inter-Variate Relationships
EEG-Specific Variate Embeddings.
In medicine, an electroencephalography (EEG) is used to record brain activity via electrodes attached to the scalp. The spatial arrangement of these electrodes is defined by established protocols, such as the international $10$-$20$ system for EEG recordings (Homan et al., 1987). Hence, in an EEG recording, each variate corresponds to the brain activity captured by a single electrode with a well-defined spatial position. We hypothesise that the spatial arrangement of these electrodes is recoverable from the EEG-specific variate embeddings, assuming that recordings from electrodes in close proximity exhibit higher correlations than those from electrodes placed farther apart. To test this hypothesis, we analyse the latent space of EEG-specific variate embeddings learned during pre-training and fine-tuning, as described in the following.
(a) $3$D spatial arrangement of EEG electrodes defined by the international $10$-$20$ system for EEG recordings.
(b) First three principal components of EEG-specific variate embeddings.
@@ E.2.1 Inter-Variate Relationships
Figure 7: Analysis of the variate embeddings for $10$-$20$ system EEG recordings with $19$ electrodes learned during pre-training. The label of each coordinate corresponds to the electrode name.
Assume 3D electrode coordinates of the international $10$-$20$ system for EEG recordings (Homan et al., 1987) to be defined in Euclidean space $\mathbb{E}_{Y}^{3}$ (see Figure 7(a)). We first determine the first three principal components of the EEG-specific variate embeddings, thus projecting them into a Euclidean space $\mathbb{E}_{X}^{3}$ (see Figure 7(b)). Then, we perform a multivariate linear regression
$\bm{Y}=\bm{1}\beta_{0}+\bm{X}\bm{B}+\bm{\epsilon}\in\mathbb{R}^{N\times 3}\quad\text{with}\quad\beta_{0}\in\mathbb{R}^{1\times 3},\bm{X}\in\mathbb{R}^{N\times 3},\bm{B}\in\mathbb{R}^{3\times 3},\bm{\epsilon}\in\mathbb{R}^{N\times 3},$ (8)
where $\bm{1}\in\mathbb{R}^{N\times 1}$ is a vector of ones and $N$ denotes the number of electrodes, to align the first three principal components in $\mathbb{E}_{X}^{3}$ (here, $\bm{X}$) with the $3$D electrode coordinates in $\mathbb{E}_{Y}^{3}$ (here, $\bm{Y}$). Finally, to quantify this very alignment, we determine the coefficient of determination $R^{2}\in[0,1]$. Note that $R^{2}=1$ represents a perfect alignment, i.e. $\mathbb{E}_{Y}^{3}=\mathbb{E}_{X}^{3}$, where the first three principal components of the EEG-specific variate embeddings only need to be shifted and scaled to recover the spatial arrangement of the electrodes (i.e. $\epsilon$ is a zero matrix).
We analyse (i) variate embeddings of $10$-$20$ system EEG recordings with $19$ electrodes learned during pre-training and (ii) variate embeddings of previously unseen $10$-$20$ system EEG recordings with $32$ electrodes learned during fine-tuning. We find that the first three principal components explain $74.7\,\%$ of the variance in (i) and $87.9\,\%$ in (ii). Moreover, we observe $R^{2}$ values of $0.81$ for (i) and $0.95$ for (ii), indicating an accurate alignment of EEG-specific variate embeddings with the electrode layout. A $2$D visualisation of this alignment is provided in Figure 8. Overall, these findings confirm our initial hypothesis that the spatial arrangement of EEG electrodes can be precisely recovered from EEG-specific variate embeddings.
(a) Variate embeddings of $10$-$20$ system EEG recordings with $19$ electrodes learned during pre-training.
(b) Variate embeddings of previously unseen $10$-$20$ system EEG recordings with $32$ electrodes learned during fine-tuning.
Figure 8: First two principal components of EEG-specific variate embeddings, overlaid on the true EEG electrode layout. The embeddings accurately reflect the spatial arrangement of the EEG electrodes, confirmed by high correlations (R2) between the principal components $\bullet$ and the electrode arrangement $\circ$.
ECG-Specific Variate Embeddings.
Another application in medicine is electrocardiography (ECG), which is widely used to measure the electrical activity of the heart via electrodes placed on the human torso. The standard $12$-lead ECG is composed of $3$ limb leads (leads I, II, and III), $3$ augmented limb leads (leads aVR, aVL, and aVF), and $6$ precordial leads (leads V1 to V6). The limb leads (or Einthoven leads Einthoven et al. (1902)) reflect the potential differences between pairs of electrodes attached to the limbs (right arm: RA, left arm: LA, and left leg: LL): I, II, and III represent the potential difference between RA and LA, RA and LL, and LA and LL, respectively. Together, these limb leads form the so called Einthoven triangle Kligfield et al. (2007), spanned by the three vectors RA $\rightarrow$ LA, RA $\rightarrow$ LL, and LA $\rightarrow$ LL. Similarly, the augmented limb leads (or Goldberger leads Goldberger (1942)) also reflect potential differences between pairs of electrodes attached to the limbs. In contrast to the previous leads, however, the precordial leads (or Wilson leads Wilson et al. (1934)) are directly measured from the electrodes placed on the chest.
When analysing the latent space of the ECG-specific variate embeddings learned during pre-training, we can clearly observe these relationships among the $12$ leads, as depicted in Figure 9. For instance, the limb leads I, II, and III form a planar $2$D triangle, thus successfully reflecting Einthoven’s triangle Kligfield et al. (2007). Additionally, the precordial leads V1111 to V6666, arranged on the rib cage from the sternum to the mid-axillary line, represent the $3$D view of the human heart.